OpenStudy (anonymous):

simplify 14-7x over x-2

OpenStudy (anonymous):

@itsmylife help?(:

OpenStudy (anonymous):

\[\frac{ 14-7x }{ x-2 }=\frac{ 7*2-7x }{ x-2 }=\frac{ 7(2-x) }{ x-2 }=\frac{ -7(x-2) }{ x-2 }=-7\]

OpenStudy (anonymous):

thanks youre awesome (:

OpenStudy (anonymous):

so are you babe ;)